1. Find a power series representation centered at x = 0 of the function x2 / (4-x2) and determine its interval of convergence.
  2. Determine all numbers x so that the series "\displaystyle\sum_{k=0}^∞" (ln x)k converges
  3. For which p is the series "\displaystyle\sum_{n=0}^∞" n/ (n2+1)p convergent?

Thank you so much.


1
Expert's answer
2021-11-10T11:49:34-0500

1. We are given f(x)="\frac{x^2}{4-x^2}"

We know "\displaystyle\sum_{n=0}^{\infty}x^n=\frac{1}{1-x} \ \ ,|x|<1" this is equation (i), Where 1 is the first term and x is the common ratio.

Here f(x)="\frac{x^2}{4-x^2}=\frac{\frac{x^2}{4}}{1-\frac{x^2}{4}}" Here first term is x2 and common ratio is (x2-3)

Relating this with equation (i)

"\frac{x^2}{4-x^2}=\frac{x^2}{4}\displaystyle\sum_{n=0}^{\infty}(\frac{x^2}{4})^n, and \ |\frac{x^2}{4}|<1"

"=\displaystyle\sum_{n=0}^{\infty}\frac{x^{2n}}{4^{2n}}\frac{x^2}{4}"

"=\displaystyle\sum_{n=0}^{\infty}\frac{x^{2(n+1)}}{4^{(2n+1)}}"

And for interval of convergence,

"|\frac{x^2}{4}|<1\implies |x^2|<4"

"\implies x^2<4"

"-2<x<2"

"\therefore" the interval of convergence is (-2,2) and the series is "\displaystyle\sum_{n=0}^{\infty}\frac{x^{2(n+1)}}{4^{(2n+1)}}"

2. The given series is "\displaystyle\sum_{k=0}^{\infty}(ln\ x)^k"

By D'Alembert's ratio test

"\lim\limits_{k\rarr\infty}\frac{a_{k+1}}{a_k}=\lim\limits_{k\rarr\infty}\frac{(ln\ x)^{k+1}}{(ln\ x)^k}"

"=\lim\limits_{k\rarr\infty}(ln\ x)^{k+1-k}"

"=\lim\limits_{k\rarr\infty}(ln\ x)"

"=ln\ x"

Since it converges, "\therefore ln\ x<1"

"\implies\ 0<x<e"

The values of x that satisfies the series is "0<x<e"

3. "\displaystyle\sum_{n=0}^{\infty}\frac{n}{(n^2+1)^p}" "\implies" Consider the series "\displaystyle\sum_{n=0}^{\infty}\frac{1}{(n^{2p-1})}"

Based on p series test

If 2p-1>1 then the series converges

2p>1+1=2

p>1

The series is convergent for p>1